Under the same condition of pressure Hydrogen diffuses 8 times as fast as gas Y. Calculate the relative molecular mass


1
Expert's answer
2022-02-23T00:44:01-0500

rate for y = 1

rate for H2 = 8


"\frac{rH_2}{rY}=\sqrt{\frac{MMY}{MMH_2}}"


"\frac{8}{1}=\sqrt\frac{MMY}{MMH_2}"


"8^2=\frac{MMY}{2}"


"MMY=64\u00d72=128g\/mol"

Relative molecular mass= 128g/mol